OpenStudy (anonymous):

PLEASE HELPPP what is the sum of the finite arithmetic series? (-15) + 0 + 15 + 30 + ... + 195

OpenStudy (anonymous):

do you want me to throw a calculator at you?

OpenStudy (anonymous):

lol I don't know how to do it

OpenStudy (anonymous):

can u tell me how?? @sourwing

OpenStudy (nikato):

This might help http://www.regentsprep.org/Regents/math/algtrig/ATP2/ArithSeq.htm

OpenStudy (anonymous):

I'M POOPING!

OpenStudy (anonymous):

@nikato thank you! but how do I fin the sum? @sourwing lol have fun

OpenStudy (anonymous):

btw I know it says how there , im looking at it

OpenStudy (nikato):

You should use An= a1 + (n-1)d to find which nth term 195 is. Plug in An=195 a1= the first term which is -15 And d=common difference which is15 And now you want to find n

OpenStudy (anonymous):

okay I got 1355

OpenStudy (anonymous):

ok, I'm done pooping. Do you still need help?

OpenStudy (nikato):

I got a different answer. We're you able to solve for n with what I told you

OpenStudy (nikato):

Well if you did it correctly. 195=-15+ (n-1)15 195=-15+15n-15 Combine like terms 195=-30+15n 225=15n n= 15

OpenStudy (nikato):

Plug this into the second equation n(a1+An) ----------- 2 15(-15+195) ------------- = ? 2

OpenStudy (anonymous):

okay now I got 3000

OpenStudy (anonymous):

@sourwing lol yes!!

OpenStudy (anonymous):

nope I got 1350! I think its right .. thaaaank you! :D

OpenStudy (anonymous):

looks like you got it